Multiply
(srt5-2sqrt6)(2sqrt5+3sqrt6)

Answer by stanbon(75887)

Multiply
(srt5-2sqrt6)(2sqrt5+3sqrt6)
------
= 2*5+6sqrt(30)-4sqrt(30)-6*6
----
= 10-36 + 2sqrt(30)
---
= -26 + 2sqrt(30)
=======================
Cheers,
Stan H.
============